A bicyclist rode into the country for 5 hours. In returning, her speed was 5 mi/hour faster and the trip took 4 hours. What was her speed each way?
r*t=d

Answer by rmromero(383)


A bicyclist rode into the country for 5 hours. In returning, her speed was 5 mi/hour faster and the trip took 4 hours. What was her speed each way?
r*t=d

What is asked in the problem?
What was her speed each way

Given:
A bicyclist rode in to the country for 5 hrs.
Returning, her speed was 5mi/h faster and the trip took 4 hrs.
Representation:
Let x = her speed on her way to the country
x + 5 = her speed when she returns

r * t = d
on the way x 5 = 5x
Returning x+5 4 = 4(x+5)
Equation:
5x = 4(x+5)
5x = 4x + 20
x = 20 mi/h her speed on her way to the country
x+5 = 20+5
= 25 mi/h her speed returning to the country



Visit my website:
http://onlinemathtutorial.tripod.com
Sign my guestbook
Thanks